# Hybrid Autoregressive-Diffusion Model for Real-Time Streaming Sign Language Production

### v1

@@ Abstract
Earlier Sign Language Production (SLP) models typically relied on autoregressive methods that generate output tokens one by one, which inherently provide temporal alignment. Although techniques like Teacher Forcing can prevent model collapse during training, they still cannot solve the problem of error accumulation during inference, since ground truth is unavailable at that stage. In contrast, more recent approaches based on diffusion models leverage step-by-step denoising to enable high-quality generation. However, the iterative nature of these models and the requirement to denoise entire sequences limit their applicability in real-time tasks like SLP. To address it, we apply a hybrid approach combining autoregressive and diffusion models to SLP for the first time, leveraging the strengths of both models in sequential dependency modeling and output refinement. To capture fine-grained body movements, we design a Multi-Scale Pose Representation module that separately extracts detailed features from distinct articulators and integrates them via a Multi-Scale Fusion module. Furthermore, we introduce a Confidence-Aware Causal Attention mechanism that utilizes joint-level confidence scores to dynamically guide the pose generation process, improving accuracy and robustness. Extensive experiments on the PHOENIX14T and How2Sign datasets demonstrate the effectiveness of our method in both generation quality and real-time streaming efficiency. Project page is available at: https://hybridsign.github.io/
@@ Introduction
Sign language serves as a vital communication medium for the deaf and hard-of-hearing communities, relying heavily on precise and expressive body and hand movements and facial expressions. The automatic generation of sign language poses significant challenges due to the complexity and subtlety of human gestures, the high-dimensional nature of pose sequences, and the need for real-time performance in practical applications. Existing approaches typically employ either autoregressive models atmc ; ptslp ; gsen , which excel at capturing temporal dependencies but may suffer from error accumulation, or diffusion models Huang2021TowardsFA ; g2pddm ; gcdm , which generate high-quality samples but are often computationally intensive and less suited for streaming scenarios.
In this work, inspired by the concept of self-forcing selfforing , we design a Hybrid Autoregressive-Diffusion Model for SLP that combines the strengths of both paradigms to enable real-time streaming generation with high fidelity and temporal coherence. Specifically, we do not adopt the teacher forcing strategy during training—instead, the model relies on its own predictions rather than ground truth data, thereby avoiding the inconsistency between training and inference.
@@
Our extensive experiments on the PHOENIX14T phoenix and How2Sign how2sign sign language datasets demonstrate that the proposed model achieves the effectiveness in terms of generation quality, temporal consistency, and computational efficiency. Moreover, the streaming capability of our hybrid model makes it highly practical for real-world sign language communication systems, including live interpretation and interactive applications.
@@
Although both autoregressive and diffusion-based models have achieved significant progress, real-time sign language production remains highly challenging. Autoregressive models naturally excel at temporally modeling and can output each frame immediately without waiting for the entire sequence to be generated, making them highly suitable for real-time scenarios. However, even with strategies like Teacher Forcing to mitigate training collapse, these models still suffer from exposure bias during inference due to their reliance on predicted frames rather than ground truth. On the other hand, diffusion models produce high-quality generation but rely on an iterative denoising process that introduces substantial latency, limiting their applicability in real-time settings. To overcome these limitations, we for the first time apply a hybrid model to SLP that leverages the advantages of both approaches. Our method is able to generate the first frame of the video in 1.22 seconds and then stream subsequent high-quality frames continuously. Compared to the methods we are aware of, it takes significantly less time and effectively addresses the challenges of real-time sign language production.
@@
Autoregressive (AR) models generate a sequence by modeling each element conditioned on its previous elements. For a sequence of human poses $P=\{P_{1},P_{2},\ldots,P_{T}\}$, where $P_{t}\in R^{J\times D}$ represents the positions of $J$ joints at time $t$ in $D$-dimensional space, the autoregressive factorization is:
@@
To preserve temporal causality in streaming generation, we apply a causal mask to prevent attention to future time steps.
@@
Our proposed framework for real-time streaming sign language production integrates three key components:
@@
As shown in Figure 1, given a natural language sentence (e.g., in German or English), the model generates a temporally consistent sequence of human poses corresponding to sign language gestures in a streaming, frame-by-frame manner. First, the Multi-Scale Pose Representation module uses dedicated Expert modules to generate three articulators: face, body, and hands, through a diffusion-based process. To support real-time generation of full motion frames, these articulators are fused into a complete pose frame by the Multi-Scale Fusion module. The fused pose is then decomposed back into the three articulators. Each part serves as a conditional input for generating the corresponding part in the next frame, following an autoregressive process. The entire generation pipeline is guided by joint confidence scores to enhance both the accuracy and realism of the output.
@@
$\text{with}\quad\tilde{C}_{<t}$ =$=\{\tilde{C}_{1},\dots,\tilde{C}_{t-1}\},\quad\tilde{C}_{t}=\{\tilde{F}_{t},\tilde{H}_{t},\tilde{B}_{t}\}$
@@
The model is trained using three loss terms:
Joint Loss.
To ensure accurate joint positions:
$\mathcal{L}_{joint}=\frac{1}{J}\sum_{j=1}^{J}|p_{j}-\hat{p}_{j}|$ (18)
where $p_{j}$ and $\hat{p}_{j}$ are ground truth and predicted joint positions.
Bone Loss.
To preserve bone orientations:
$\mathcal{L}_{bone}=\frac{1}{B}\sum_{b=1}^{B}(q_{b}-\hat{q}_{b})^{2}$ (19)
where $q_{b}$ and $\hat{q}_{b}$ denote ground truth and predicted bone orientations.
Soft-DTW Loss.
To align pose sequences temporally, we use Soft-DTW softdtw :
$\mathcal{L}_{\text{soft-dtw}}=\min\nolimits_{A}^{\gamma}\sum_{(i,j)\in A}\|x_{i}-y_{j}\|^{2}$
with the smooth minimum operator:
$\min\nolimits^{\gamma}\{a_{1},\dots,a_{k}\}=-\gamma\log\left(\sum_{i=1}^{k}e^{-a_{i}/\gamma}\right)$
Total Loss.
Dynamic weighting based on inverse EMA dwa balances all terms:
$\lambda_{i}^{(t)}=\frac{1}{\bar{\mathcal{L}}_{i}^{(t)}+\epsilon}\bigg{/}\sum_{j=1}^{N}\frac{1}{\bar{\mathcal{L}}_{j}^{(t)}+\epsilon}$ (20)
The final total loss is:
$\mathcal{L}_{\text{total}}^{(t)}=\lambda_{1}^{(t)}\mathcal{L}_{joint}^{(t)}+\lambda_{2}^{(t)}\mathcal{L}_{bone}^{(t)}+\lambda_{3}^{(t)}\mathcal{L}_{soft\text{-}dtw}^{(t)}$ (21)
This scheme balances the losses to stabilize training.
@@
The benchmark datasets we tested in our experiments are as follows:
  • The RWTH-PHOENIX-Weather 2014T dataset phoenix is a German Sign Language (DGS) corpus that consists of 8,257 sentence-level samples, encompassing 2,887 unique German words. Specifically, it includes 7,096 training samples (featuring 1,066 distinct sign glosses in the gloss annotations and 2,887 words in the corresponding German spoken language translations), 519 validation samples, and 642 test samples.
  • The How2Sign dataset how2sign is a large-scale American Sign Language (ASL) corpus designed for sign language understanding and generation tasks. It contains over 80 hours of multimodal data, including RGB video and 2D keypoints. The dataset training dataset consists of 31,048 video segments, with 1,739 segments in the development set and 2,343 segments in the test set.
@@
Following the existing works g2pddm ; gcdm ; signidd ; signd2c , we also employed four types of metrics to evaluate the model’s performance, including BLEU bleu , ROUGE rouge , WER, DTW and FID. However, since SLP cannot directly compute these metrics, we employed a pre-trained SLT model slt for back-translation. Specifically, the generated sign language videos are translated back into natural language, and the original text input is compared with the translated output to calculate these metrics.
@@
Table 3: Latency and throughput comparison between the baseline model and this model with different numbers of expert modules for generating a 5-second, 60-frame sign language pose sequence. The total time includes processing by the text encoder and the generative model.
Method Latency (s)$\downarrow$ Throughput (FPS)$\uparrow$
@@
HybridSign with 4 experts 2.02 5.66
HybridSign with 3 experts 1.22 10.17
  • Note: ”4 experts” refer to face, body, left hand, and right hand. ”3 experts” refer to face, body, and hands combined.
@@
Table 3 compares latency and throughput with the baseline model. Here, latency refers to the time required to generate the first frame of the 60-frame sequence, rather than the total time to produce the entire video. This reflects the model’s initial response speed, which is important for real-time applications. HybridSign achieves significantly lower latency and higher throughput than the baseline, indicating its efficiency and practicality in real-time sign language production.
Performance differences across the two datasets are relatively small, likely due to the use of 2D pose sequences. While this may slightly reduce generation quality, it lowers computational complexity and improves training and inference efficiency. To compensate, we incorporate the Multi-Scale Pose Representation Module and the Confidence-Attention Causal Mechanism, which enhance overall performance. Notably, the DTW score is reduced by approximately 80% thanks to the Soft-DTW loss softdtw , which improves temporal alignment.
@@
Table 4: Ablation study results of modules on the How2Sign dataset.
Modules DEV TEST
@@
RNN 23.47 5.02 6.89 22.92 4.72 7.84
Causal Attention 26.46 5.71 5.10 25.08 5.83 5.50
Confidence-Aware Causal Attention 30.71 6.92 3.76 30.12 6.48 3.89
@@
1 (whole pose) 22.73 5.31 6.85 22.33 5.14 7.02
4 (face + body + lh + rh) 30.83 6.98 3.72 30.31 6.56 3.77
3 (face + body + hands) 30.71 6.92 3.76 30.12 6.48 3.89
@@
We also conducted two ablation studies to evaluate the effects of using different attention mechanisms and varying numbers of expert modules.
Table 4 shows the performance differences among various autoregressive implementations. The model using Confidence-Aware Causal Attention outperforms those using RNN rnn or standard Causal Attention attention . This is because the attention mechanism takes confidence scores into account, assigning different attention weights to keypoints based on their reliability.
Table 5 presents the results of varying the number of expert modules. When only a single expert module is used—treating the entire pose as a whole—the generation quality drops significantly due to the inability to capture local details. Using three or four expert modules yields similar performance, but adding more experts increases the number of parameters, leading to higher computational cost 3. Therefore, we choose to use three expert modules as a balanced trade-off between performance and efficiency.
@@ Conclusions
This paper is the first to apply a hybrid framework that combines autoregressive and diffusion models to sign language production, enabling real-time, streaming generation of sign language pose frames. The framework leverages the inherent sequential modeling capability and frame-by-frame output of autoregressive models, along with the high-quality generation strengths of diffusion models. Additionally, it incorporates multi-scale pose representation modeling and a confidence-aware causal attention mechanism, which further improve the naturalness and expressive accuracy of the generated sign language. Extensive experiments on several benchmark datasets demonstrate the effectiveness of the proposed method, especially in real-time application scenarios.

### v3

# Hybrid Autoregressive-Diffusion Model for Real-Time Sign Language Production

@@ Abstract
Earlier Sign Language Production (SLP) models typically relied on autoregressive methods that generate output tokens one by one, which inherently provide temporal alignment. Although techniques like Teacher Forcing can prevent model collapse during training, they still cannot solve the problem of error accumulation during inference, since ground truth is unavailable at that stage. In contrast, more recent approaches based on diffusion models leverage step-by-step denoising to enable high-quality generation. However, the iterative nature of these models and the requirement to denoise entire sequences limit their applicability in real-time tasks like SLP. To address it, we explore a hybrid approach that combines autoregressive and diffusion models for SLP, leveraging the strengths of both models in sequential dependency modeling and output refinement. To capture fine-grained body movements, we design a Multi-Scale Pose Representation module that separately extracts detailed features from distinct articulators and integrates them via a Multi-Scale Fusion module. Furthermore, we introduce a Confidence-Aware Causal Attention mechanism that utilizes joint-level confidence scores to dynamically guide the pose generation process, improving accuracy and robustness. Extensive experiments on the PHOENIX14T and How2Sign datasets demonstrate the effectiveness of our method in both generation quality and real-time efficiency.
@@ Introduction
Sign language serves as a vital communication medium for the deaf and hard-of-hearing communities, relying heavily on precise and expressive body and hand movements and facial expressions. The automatic generation of sign language poses significant challenges due to the complexity and subtlety of human gestures, the high-dimensional nature of pose sequences, and the need for real-time performance in practical applications. Existing approaches typically employ either autoregressive models atmc ; ptslp ; gsen , which excel at capturing temporal dependencies but may suffer from error accumulation, or diffusion models Huang2021TowardsFA ; g2pddm ; gcdm , which generate high-quality samples but are often computationally intensive and less suited for real-time scenarios.
In this work, inspired by the concept of self-forcing selfforing , we design a Hybrid Autoregressive-Diffusion Model for SLP that combines the strengths of both paradigms to enable real-time generation with high fidelity and temporal coherence. Specifically, we do not adopt the teacher forcing strategy during training—instead, the model relies on its own predictions rather than ground truth data, thereby avoiding the inconsistency between training and inference.
@@
Our extensive experiments on the PHOENIX14T phoenix and How2Sign how2sign datasets demonstrate that the proposed model delivers high generation quality, strong temporal consistency, and notable computational efficiency, making it highly practical for real-world sign language communication systems, including live interpretation and interactive applications.
@@
Although both autoregressive and diffusion-based models have achieved significant progress, real-time sign language production remains highly challenging. Autoregressive models naturally excel at temporally modeling and can output each frame immediately without waiting for the entire sequence to be generated, making them highly suitable for real-time scenarios. However, even with strategies like Teacher Forcing to mitigate training collapse, these models still suffer from exposure bias during inference due to their reliance on predicted frames rather than ground truth. On the other hand, diffusion models produce high-quality generation but rely on an iterative denoising process that introduces substantial latency, limiting their applicability in real-time settings. To overcome these limitations, we for the first time apply a hybrid model to SLP that leverages the advantages of both approaches. Our method is able to generate the first frame of the video in 5.9 seconds and then subsequent high-quality frames continuously. Compared to the methods we are aware of, it takes significantly less time and effectively addresses the challenges of real-time sign language production.
@@
Autoregressive models generate a sequence by modeling each element conditioned on its previous elements. For a sequence of human poses $P=\{P_{1},P_{2},\ldots,P_{T}\}$, where $P_{t}\in R^{J\times D}$ represents the positions of $J$ joints at time $t$ in $D$-dimensional space, the autoregressive factorization is:
@@
To preserve temporal causality in generation, we apply a causal mask to prevent attention to future time steps.
@@
Our proposed framework for real-time sign language production integrates three key components:
@@
As shown in Figure 1, given a natural language sentence (e.g., in German or English), the model generates a temporally consistent sequence of human poses corresponding to sign language gestures in a frame-by-frame manner. First, the Multi-Scale Pose Representation module uses dedicated Expert modules to generate three articulators: face, body, and hands, through a diffusion-based process. To support real-time generation of full motion frames, these articulators are fused into a complete pose frame by the Multi-Scale Fusion module. The fused pose is then decomposed back into the three articulators. Each part serves as a conditional input for generating the corresponding part in the next frame, following an autoregressive process. The entire generation pipeline is guided by joint confidence scores to enhance both the accuracy and realism of the output.
@@
$\text{with}\quad\tilde{C}_{<t}$ $=\{\tilde{C}_{1},\dots,\tilde{C}_{t-1}\},\quad\tilde{C}_{t}=\{\tilde{F}_{t},\tilde{H}_{t},\tilde{B}_{t}\}$
@@
Our model is optimized using a composite loss that combines three complementary objectives: joint accuracy, kinematic consistency, and temporal alignment.
Joint loss
To encourage accurate joint localization, we minimize the mean absolute error between predictions and ground truth:
$\mathcal{L}_{\mathrm{joint}}\;=\;\frac{1}{J}\sum_{j=1}^{J}\|p_{j}-\hat{p}_{j}\|_{1},$ (18)
where $p_{j}\in\mathbb{R}^{d}$ and $\hat{p}_{j}\in\mathbb{R}^{d}$ denote the ground-truth and predicted position of joint $j$, and $\|\cdot\|_{1}$ is the $\ell_{1}$ norm.
Bone loss
To preserve bone orientations and kinematic consistency, we penalize orientation deviations:
$\mathcal{L}_{\mathrm{bone}}\;=\;\frac{1}{B}\sum_{b=1}^{B}\|q_{b}-\hat{q}_{b}\|_{2}^{2},$ (19)
where $q_{b}$ and $\hat{q}_{b}$ are the ground-truth and predicted orientation (e.g. unit quaternions or axis-angle vectors) of bone $b$, and $\|\cdot\|_{2}$ denotes the Euclidean norm. If orientations are represented as quaternions, consider using a geodesic/angular metric as an alternative.
Soft-DTW loss
To align predicted and ground-truth pose sequences temporally, we employ differentiable Dynamic Time Warping (Soft-DTW) softdtw :
$\mathcal{L}_{\mathrm{soft\text{-}dtw}}\;=\;\min\nolimits_{A}^{\gamma}\sum_{(i,j)\in A}\|x_{i}-y_{j}\|_{2}^{2},$ (20)
where $x_{i}$ and $y_{j}$ are per-frame pose descriptors, $A$ is an alignment path, and $\gamma>0$ controls the smoothness of the soft minimum:
$\min\nolimits^{\gamma}\{a_{1},\dots,a_{k}\}=-\gamma\log\!\Bigg(\sum_{i=1}^{k}e^{-a_{i}/\gamma}\Bigg).$ (21)
Adaptive total loss
We balance the three terms with dynamic weights computed from the inverse exponential moving average (EMA) of each loss dwa . Let $\bar{\mathcal{L}}_{i}^{(t)}$ be the EMA of loss $i$ at iteration $t$, and $\epsilon>0$ a small constant to avoid division by zero. The weight for loss $i$ is
$\lambda_{i}^{(t)}\;=\;\frac{\big(\bar{\mathcal{L}}_{i}^{(t)}+\epsilon\big)^{-1}}{\sum_{j=1}^{N}\big(\bar{\mathcal{L}}_{j}^{(t)}+\epsilon\big)^{-1}},\qquad i=1,\dots,N,$ (22)
with $N=3$ in our case. The final objective at iteration $t$ is
$\mathcal{L}_{\mathrm{total}}^{(t)}\;=\;\lambda_{1}^{(t)}\,\mathcal{L}_{\mathrm{joint}}^{(t)}\;+\;\lambda_{2}^{(t)}\,\mathcal{L}_{\mathrm{bone}}^{(t)}\;+\;\lambda_{3}^{(t)}\,\mathcal{L}_{\mathrm{soft\text{-}dtw}}^{(t)}.$ (23)
This adaptive weighting scheme emphasizes losses that are relatively smaller (via inverse EMA), promoting stable and balanced training across spatial accuracy, kinematic consistency, and temporal alignment.
@@
We evaluate the proposed method on two benchmarks: PHOENIX14T and How2Sign. The PHOENIX14T phoenix is a German Sign Language (DGS) corpus that consists of 8,257 sentence-level samples, encompassing 2,887 unique German words. The How2Sign how2sign is a large-scale American Sign Language (ASL) corpus designed for sign language understanding and generation tasks. It contains over 80 hours of multimodal data.
@@
Following the existing works g2pddm ; gcdm ; signidd ; signd2c , we used a pre-trained SLT model slt for back-translation, converting generated sign language pose sequences back to text, and then evaluated the results with BLEU bleu , ROUGE rouge , WER, DTW, and FID.
@@
Table 3: Latency and throughput comparison between the baseline models and this model with 3 expert modules for generating 60-frame sign language pose sequence.
[…]
GCDM gcdm 52.18 1.15
@@
G2P-DDM g2pddm 25.78 2.33
HybridSign 5.90 10.17
@@
Table 3 compares latency and throughput with the baseline models. Here, latency refers to the time required to generate the first frame of the 60-frame sequence, rather than the total time to produce the entire video. This reflects the model’s initial response speed, which is important for real-time applications. HybridSign achieves significantly lower latency and higher throughput than the baselines, indicating its efficiency and practicality in real-time sign language production.
Performance differences across the two datasets are relatively small, likely due to the use of 2D pose sequences. While this may slightly reduce generation quality, it lowers computational complexity and improves training and inference efficiency. To compensate, we incorporate the Multi-Scale Pose Representation Module and the Confidence-Attention Causal Mechanism, which enhance overall performance. Notably, the DTW score is reduced by approximately 20% thanks to the Soft-DTW loss softdtw , which improves temporal alignment.
@@
Table 4: Ablation study results of modes on the How2Sign dataset.
@@
Diffusion Mode 30.25 6.55 8.06 32.89 1.83
Autoregressive Mode 26.15 5.40 4.49 5.53 10.85
Hybrid Mode 30.12 6.48 3.89 5.90 10.17
Table 5: Ablation study results of modules on the How2Sign dataset. CA denotes Causal Attention, CACA denotes Confidence-Aware Causal Attention.
[…]
RNN 23.47 5.02 6.89 4.72 9.71
CA 25.08 5.83 5.50 5.48 10.95
CACA 30.12 6.48 3.89 5.90 10.17
@@
1 (whole pose) 22.33 5.14 7.02 7.69 7.80
4 (face + body + lh + rh) 29.17 6.03 5.72 7.73 7.76
3 (face + body + hands) 30.12 6.48 3.89 5.90 10.17
@@
We conducted three ablation studies on the generation of 60-frame sign language pose sequences to systematically evaluate our method. Specifically, we compared the proposed hybrid approach with a pure diffusion model and a pure autoregressive model, and further investigated the impact of different attention mechanisms and varying numbers of expert modules.
Table  4 compares the performance and efficiency of a pure diffusion model and a pure autoregressive model. Specifically, the diffusion model adopts the DiT architecture  dit , while the autoregressive model uses a Transformer decoder with teacher forcing  attention . Experimental results show that, thanks to its multi-step denoising process, the diffusion model achieves very high generation quality and reaches the best scores on both the B1 and B4 metrics. However, its temporal consistency and generation efficiency lag significantly behind the other models, making it unsuitable for real-time scenarios. In contrast, the autoregressive model generates frame-by-frame, resulting in extremely low first-frame latency, but suffers from a noticeable quality drop due to the mismatch between training and inference. Our proposed approach combines the high-quality generation of diffusion models with the frame-by-frame generation of autoregressive models, achieving both strong performance and efficiency.
Table 5 presents the performance differences among various autoregressive implementations, including a recurrent neural network (RNN) rnn , a standard causal attention mechanism attention , and a confidence-aware causal attention mechanism. The results indicate that the RNN-based model achieves very low generation quality, but due to its computational simplicity compared to attention mechanisms, it exhibits excellent first-frame latency. However, its strict sequential dependency limits overall efficiency, resulting in lower throughput than attention-based methods. In contrast, the confidence-aware causal attention mechanism incorporates confidence scores, assigning different attention weights based on the reliability of keypoints, which leads to the best generation quality while maintaining strong applicability in real-time scenarios.
Table 6 presents the results of varying the number of expert modules. The experiments include using a single expert module (treating the entire pose as a whole), using three expert modules to handle the face, body, and hands separately, and using four expert modules to handle the face, body, left hand, and right hand individually. The results show that with a single expert, the model cannot capture local details, leading to a significant drop in generation quality. Moreover, since the single module cannot take advantage of parallel computation, its efficiency is relatively low. Using four experts increases inter-module dependencies and the difficulty of feature fusion, which can cause information conflicts or redundancy and negatively affect generation quality. Although parallelization eliminates most of the serial computation overhead, the integration of outputs from multiple modules still requires some serial processing, limiting overall performance. In contrast, using three expert modules strikes the optimal balance, achieving the best trade-off between generation quality and efficiency.
@@ Conclusions
This paper explores a hybrid framework that combines autoregressive and diffusion models for sign language production, enabling real-time generation of individual sign language pose frames. The framework leverages the inherent sequential modeling capability and frame-by-frame output of autoregressive models, along with the high-quality generation strengths of diffusion models. Additionally, it incorporates multi-scale pose representation modeling and a confidence-aware causal attention mechanism, which further improve the naturalness and expressive accuracy of the generated sign language. Extensive experiments on several benchmark datasets demonstrate the effectiveness of the proposed method, particularly in real-time applications.